OpenStudy (anonymous):

simplify 5^3sqrt96 simplify 5^3sqrt96 @Mathematics

OpenStudy (anonymous):

\[5\sqrt[3]{96}\]

OpenStudy (anonymous):

96 = 2 x 2 x 2 x 2 x 2 x 3 96 = 2^5 x 3 cubic root ( 2^5 x 3) = 2 cubic root(2^2 x 3) answer is \[10\sqrt[3]{12}\]

myininaya (myininaya):

\[5 \cdot \sqrt[3]{8 \cdot 12}=5 \cdot \sqrt[3]{2^3 \cdot 12}=5 \cdot \sqrt[3]{2^3} \cdot \sqrt[3]{12}\]

myininaya (myininaya):

\[5(2)\cdot \sqrt[3]{12} =10 \cdot \sqrt[3]{12}\]